a particle is projected
vertically upward from a point A on the ground . It takes t1 time to reach the point  B but it still continues to move up . if it take further t2 time to reach the ground from point B then the hight of the point B from the ground will be
 
  1. 1/2g (t1+t2)2
  2. gt1t2
  3. 1/8g(t1+t2)2
  4. 1/2gt1+t2

Grade:11

1 Answers

Arun
4 years ago
Time taken for the particle to reach highest point is t1+t22t1+t22
(v=u−gt)(v=u−gt)
Therefore initial velocity=u=g(t1+t22)=u=g(t1+t22)
Therefore heigh of B from ground is
h=ut+12h=ut+12at2at2
h=ut1−12h=ut1−12gt12gt12
=g(t1+t22)=g(t1+t22)t1−12t1−12gt21gt12
h=12h=12gt1t2